### v1

# Event-Based Tracking Any Point with Motion-Augmented Temporal Consistency

Abstract
Tracking Any Point (TAP) plays a crucial role in motion analysis. Video-based approaches rely on iterative local matching for tracking, but they assume linear motion during the blind time between frames, which leads to target point loss under large displacements or nonlinear motion. The high temporal resolution and motion blur-free characteristics of event cameras provide continuous, fine-grained motion information, capturing subtle variations with microsecond precision. This paper presents an event-based framework for tracking any point, which tackles the challenges posed by spatial sparsity and motion sensitivity in events through two tailored modules. Specifically, to resolve ambiguities caused by event sparsity, a motion-guidance module incorporates kinematic features into the local matching process. Additionally, a variable motion aware module is integrated to ensure temporally consistent responses that are insensitive to varying velocities, thereby enhancing matching precision. To validate the effectiveness of the approach, an event dataset for tracking any point is constructed by simulation, and is applied in experiments together with two real-world datasets. The experimental results show that the proposed method outperforms existing SOTA methods. Moreover, it achieves 150% faster processing with competitive model parameters. The project page is here.
1 Introduction
Figure 1: Video-based point tracking method (first row) face limitations in tracking objects with varying motion states, primarily due to their reliance on slow linear motion assumptions during blind times. Our approach (second row) leverages continuous motion information from events, achieving smoother and more accurate results.
Tracking Any Point (TAP) aims to determine the subsequent positions of a given query point on a physical surface over time, which is essential for understanding object motion in the scene. It becomes even more vital for autonomous driving and embodied agents [37, 6, 31], where operations require precise spatial control of objects over time.
(a) The spatio-temporal properties of events.
(b) The solution of this paper.
Figure 2: (a) The spatio-temporal distribution of events generated by a stick rotating uniformly around a pivot. Sampling from different patches along the stick reveals spatial sparsity of events. Counting the events at each patch shows a positive correlation between event update frequency and object speed, which causes temporal inconsistencies with varying motion speeds. (b) This paper leverages the temporal continuity of events to capture subtle motion changes, enhancing the matching process and thereby improving tracking accuracy.
Recent methods rely on video input, predicting the positions of query points by matching their appearance features with local regions in subsequent frames [9, 10, 15]. However, as these methods assume slow linear motion during the blind time between frames, they faces the challenge of objects may undergo large displacements or nonlinear motion, causing query points to exceed the bounds of local regions and resulting in ambiguities in feature matching, see Fig. 1. While some approaches attempt to mitigate this by considering spatial context [36, 7], they still struggle to overcome the lack of motion during blind time.
To cope with the above issue, this paper utilize event cameras to capture the motion during blind time. Event cameras [22, 33] are bio-inspired sensors that respond to pixel-level brightness changes with microsecond temporal resolution, generating sparse and asynchronous event streams. They have the characteristics of high temporal resolution, no motion blur, and low energy consumption. Therefore, parsing motion during blind times using event streams is a feasible solution for tracking any point. Furthermore, capturing motion alone significantly reduces computational overhead compared to traditional frame-based cameras, enabling more efficient methods.
However, as shown in Fig. 2(a), the unique spatio-temporal properties of events make existing methods difficult to apply, manifesting in two ways: 1) Event cameras respond only to pixels with brightness changes, resulting in a sparse spatial distribution that introduces ambiguities in appearance-based matching due to spatial discontinuities. 2) The variable speed of scene objects causes fluctuations in event update frequencies, impacting the temporal consistency of event representations. Higher motion speeds lead to increased event update frequencies and densities, and vice versa. Density variations cause inconsistencies in event representations over time, affecting the precision and reliability of temporal matching.
To address these issues, this study leverages the temporal continuity of events to guide matching, alleviating ambiguities caused by spatial sparsity, as illustrated in Fig. 2(b). The temporal dynamics of events capture subtle variations in motion trajectory, complementing spatial appearance to provide coherent matching cues. Additionally, by modeling kinematic features to estimate object speed and motion patterns, the method dynamically adjusts appearance feature extraction, ensuring temporally consistent responses.
Therefore, this paper proposes a novel event-based point tracking framework that comprises a Motion-Guidance Module (MGM) and a Variable Motion Aware Module (VMAM). Specifically, MGM leverages the gradient from event time surface to compute kinematic features that construct a dynamic-appearance space, clarifying ambiguities in feature matching and guiding the extraction of temporally consistent appearance features. Additionally, to model the non-stationary states of events, VMAM is introduced to employ kinematic cues for adaptive correction and generate temporally consistent responses by combining long-term memory parameters with hierarchical appearance. The method is evaluated on a synthetic dataset as well as two real-world datasets, with experimental results demonstrating its superiority. The main contributions of this paper can be summarized as follows:
  • An event-based framework for tracking any point is presented to monitor surface points on objects by leveraging the temporal continuity of events.
  • This paper reveals the impact of spatial sparsity and motion sensitivity in event data on TAP. To address these limitations, a motion-guidance module is proposed to enhance matching process using temporal continuity, while a variable motion aware module models kinematic cues to estimate non-stationary event states, thereby improving tracking accuracy.
  • Experimental results demonstrate that the proposed approach outperforms current state-of-the-art methods on both synthetic and real-world datasets. Moreover, the proposed method exhibits competitive model parameters and 150% faster computational speed.
2 Related Work
2.1 Event-based Optical Flow
Event-based optical flow estimation can be categorized into model-based and learning-based methods. Model-based methods rely on physical priors [28, 1, 14, 30], primarily using contrast maximization to estimate optical flow by minimizing edge misalignment. Unfortunately, the strict motion assumptions inherent in these methods struggle in complex scenes, leading to decreased accuracy.
Learning-based methods have significantly improved the quality of optical flow estimation [40, 8, 13, 35, 42, 41, 34]. They can be classified into unsupervised [40, 41, 42] and supervised approaches [8, 13, 35, 34]. For the former, several methods have been proposed, including those that use event cameras alone [41, 42] or in combination with other data modalities [40]. They employ motion compensation to warp and align data across time to construct a loss function. The latter commonly utilize a coarse-to-fine pyramid structure or iterative optimization to refine the estimation. For example, Gehrig et al. [12] propose E-RAFT, which mimics iterative optimization algorithms by updating the correlation volume to optimize optical flow.
Although these methods have achieved promising results, challenges remain when applying them to TAP. Since optical flow is computed over neighboring time, linking motion vectors over long time leads to error accumulation. Additionally, optical flow calculates the correspondence between pixel points, rather than physical surface points.
Figure 3: (a) Framework overview. Given the event data and the initial positions of target points as input, the model initializes the locations for subsequent time steps, along with appearance features. It then iteratively calculates kinematic features and updates the appearance correlation map at each point to refine the trajectory. (b) Motion-Guidance Module. MGM extracts kinematic features from the gradient information in the event stream, guiding appearance feature matching and forming a dynamic-appearance matching space with the appearance features. (c) Variable Motion Aware Module. VMAM leverages kinematic features from MGM to produce temporally consistent feature responses, thereby resulting in robust correlation maps.
2.2 Event-based Feature Tracking
Feature tracking aims to predict the trajectories of keypoints. Early approaches can be grouped into two types: one [20, 39] treats feature points as event sets and tracks them using the ICP [5] method, while the other [11] extracts feature blocks from reference frames and matches them by computing brightness increments from events. Moreover, event-by-event trackers [2, 3, 18] exploit the inherent asynchronicity of event streams. Unfortunately, these methods involve complex model parameters that require extensive manual tuning for different event cameras and new environments.
To tackle these deficiencies, learning-based feature tracking methods have gained attention from researchers [25, 21]. DeepEvT [25] is the first data-driven method for event feature tracking. [21] expands 2D feature tracking to 3D and collected the first event 3D feature tracking dataset.
However, existing methods track high-contrast points by relying on local feature descriptors. TAP requires the ability to track points in low-texture areas, where current methods struggle to establish reliable descriptors.
2.3 Tracking Any Point
Tracking any point based on events has yet to be proposed, while techniques for using standard frames have been developed [9, 10, 15, 7, 38]. These methods model the appearance around the points, using MLPs to capture long-range temporal contextual relationships across frames. During inference, they employ a sliding time window to handle long videos. For example, PIPs [15] frames pixel tracking as a long-range motion estimation problem, updating trajectories through iterative local searches. In contrast, TAP-Net [9] formalizes the problem as tracking any point, overcoming occlusion through global search. However, these methods track points independently, leading to ambiguities in feature matching. Consequently, some studies [36, 7, 19] have been proposed to utilize spatial context to alleviate this issue. In addition to methodological innovations, the PointOdyssey dataset [38] is collected to advance the field, featuring the longest average video length and the highest number of tracked points to date.
Unfortunately, applying these methods to event data encounters several limitations. The spatial sparsity of events leads to misalignment when solely modeling based on appearance. Additionally, the temporal inconsistency in event density, caused by variable-speed motion, affects temporal context modeling. In this paper, a motion-guidance module is designed to construct a dynamic-appearance matching space, thereby reducing matching ambiguity. Moreover, a variable motion aware module is employed to generate temporally consistent responses for correlation operations. The method is trained and tested on simulated event modalities derived from the [38] dataset.
3 Method
3.1 Setup and Overview
Tracking any point process typically involves two stages: initializing tracking points and features, and iteratively updating point positions and associated features. Following this pipeline, an event-based method is proposed for tracking any point, as shown in Fig. 3.
Specifically, let $E_{j}=\left\{\left(x_{k},y_{k},t_{k},p_{k}\right)\right\}_{k=1}^{N}$ denote the event stream from $t_{j-1}$ to $t_{j}$, where $N$ is the number of events, and each event is a 4-tuple consisting of the coordinates $x_{k}$ and $y_{k}$, timestamp $t_{k}$, and polarity $p_{k}\in\{-1,+1\}$. Given a target point $x_{src}\in\mathbb{R}^{2}$, subsequent events are represented by the Time Surface (TS) [26], where each pixel records the timestamp of the most recent event, capturing motion process over a period of time.. During the initialization period, the TS representation is fed into a residual network [16] to extract features $\left\{F_{0},F_{1},...,F_{t}\right\}$. The point trajectories at all time steps are initialized as:
$\centering X^{0}=\left\{x_{0}^{0},x_{1}^{0},...,x_{t}^{0}\right\}=\left\{x_{src},x_{src},...,x_{src}\right\},\@add@centering$ (1)
with the corresponding feature at time $t$ being $f_{t}^{0}=F_{t}(x_{t}^{0})$.
At the iterative stage, let $x_{t}^{k}$ represent the coordinate of the query point at time $t$ after the $k$-th iteration. This paper employs VMAM to model the non-stationary states of point features at times $t-2$, $t-4$, and the initial moment, leveraging these to compute correlations $c_{t}^{k}$ with the neighboring features around $x_{t}^{k}$ at time $t$. A transformer takes correlations $C^{k}$, the kinematic features $V^{k}$ derived from MGM, along with the position-encoded apparent point motions $X_{t}^{k}-X_{t-1}^{k}$ as input to obtain the point displacement $\Delta X$. Subsequently, the point coordinates are updated through Eq. 2:
$\centering X^{k+1}=X^{k}+\Delta X,\@add@centering$ (2)
and the process is repeated iteratively.
3.2 Motion-Guidance Module
To reduce ambiguities arising from appearance-only matching, a motion-guidance module is designed to leverage the gradient characteristics of the TS to compute kinematic features, thereby building a dynamic-appearance matching space and subsequently guiding the extraction of temporally consistent appearance features.
Specifically, the event stream after TS encoding is visualized as a surface in the $xyt$ spacetime domain, representing the active events surface $\Sigma_{e}$ [4]. The spatial gradients of this surface describe the temporal changes relative to spatial variations, establishing a derivative relationship with pixel displacement at corresponding positions, see Eq. 3
$\centering\frac{\partial\Sigma_{e}}{\partial x}=\left(\frac{\partial x}{\partial\Sigma_{e}}\right)^{-1}=\left(\frac{\partial x}{\partial t}\right)^{-1}=\frac{1}{v}.\@add@centering$ (3)
For the query point $x_{t}^{k}$, this paper treats the surface formed by neighboring pixels as a surface of active events. Using the coordinates of neighboring points $(x_{1},y_{1},t_{1})$, $(x_{2},y_{2},t_{2})$, …, a system in Eq. 4 is constructed:
$\centering\begin{bmatrix}x_{1}&y_{1}&t_{1}&1\\ x_{2}&y_{2}&t_{2}&1\\ \vdots&\vdots&\vdots&\vdots\end{bmatrix}\begin{bmatrix}a\\ b\\ c\\ d\end{bmatrix}=0.\@add@centering$ (4)
Here, $(a,b,c,d)$ are coefficients for the tangent plane. The spatial gradient of the surface is estimated through plane fitting using SVD, providing kinematic vectors at the target pixel. However, in the presence of overlapping moving objects, the pixels at boundary intersections can exhibit multiple distinct motion states, disrupting local smoothness and resulting in deviations in the kinematic vectors. To address this, the proposed motion-guidance module employs a multi-layer perceptron to capture the temporal motion relationships, dynamically assigning weights to the kinematic features at different time steps, thereby correcting the motion ambiguity at the boundaries.
The corrected kinematic features serve two key roles: first, as input to VMAM, guiding the generation of speed-insensitive appearance features for correlation; second, as inputs to the transformer, combining with correlation maps and position-encoded point motions to construct a dynamic-appearance space for iterative point displacement updates.
3.3 Variable Motion Aware Module
Objects in the scene exhibit variable speeds in two primary ways: either different objects move at distinct speeds or individual objects vary their speed over time. The variation in speed affects the frequency of event updates. Assuming an object moves at speed $\boldsymbol{v}=(v_{x},v_{y})$, the illuminance change rate at any point on the edge is given by:
$\centering\frac{dI(x,y,t)}{dt}=\frac{\partial I}{\partial x}v_{x}+\frac{\partial I}{\partial y}v_{y}+\frac{\partial I}{\partial t},\@add@centering$ (5)
where $I(x,y,t)$ represents the illuminance at position $(x,y)$ at time $t$. To simplify calculations, assuming consistent global illumination, so $\frac{\partial I}{\partial t}$ is 00, and Eq. 5 simplifies to:
$\centering\frac{dI(x,y,t)}{dt}=\frac{\partial I}{\partial x}v_{x}+\frac{\partial I}{\partial y}v_{y}=\boldsymbol{v}\cdot\nabla I.\@add@centering$ (6)
Here, $\nabla I$ represents the spatial gradient of illuminance at the specified position $(x,y)$.
The event generation process can be formulated as
$log\mathcal{I}(x,y,t)-log\mathcal{I}(x,y,t-\Delta t)=pC,$ (7)
where $\Delta t$ is the time interval between consecutive events and $C$ is the contrast threshold of the event camera. This equation indicates that an event triggers when the logarithmic illuminance change at a location exceeds the threshold $C$. Combining Eqs. 6 and 7, it can be observed that
$f\propto\frac{\boldsymbol{v}\cdot\nabla I}{C},$ (8)
where $f$ signifies the event update frequency. Since $\nabla I$ depends only on the material properties of the object, $f$ is directly proportional to $\boldsymbol{v}$. In other words, higher speeds lead to higher event update frequencies, and vice versa.
Point position updates rely on consistent appearance feature matching over time. However, velocity variations disrupt the stability of event temporal distribution, causing significant matching errors. To tackle the problem, VMAM is introduced to guide appearance feature matching using kinematic features extracted through MGM. Specifically, to obtain the correlation map $c_{t}^{k}$ at the $k$-th iteration and time $t$, VMAM samples point features from the initial, $t-4$ and $t-2$ time at corresponding coordinates. These features are concatenated and fused with temporally contextual kinematic features via cross-attention. The fused features are then divided into two branches: one captures short-term temporal dependencies via 1D temporal convolution, while the other extracts long-term dependencies through temporal attention. The short-range and long-range features are summed together and correlated with the spatial context features of $x_{t}^{k}$ at time $t$, yielding $c_{t}^{k}$.
4 Experiment
Figure 4: Some examples from the Ev-PointOdyssey dataset. Each example displays the RGB image in the top-left corner, with the event modality visualization in the bottom-right corner.
Table 1: The performance of the evaluated trackers on the Ev-PointOdyssey dataset are reported in terms of $\sigma_{avg}$, MTE, $Survival_{50}$. $\sigma_{avg}$ reflects the proportion of the error between the predicted and the ground truth within a certain range. MTE represents the error between the predicted trajectory and the ground truth. $Survival_{50}$ indicates the duration of tracking. ”Dark red” and ”Orange” represent feature tracking and optical flow models. Best results are in bold; second-best are underlined.
  Methods Modality Ev-PointOdyssey Params FPS
$\sigma_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  PIPs [15] Video 0.273 0.640 0.423 28.7M 119.1
PIPs++ [38] 0.336 0.270 0.505 17.6M 122.5
TAPIR [10] 0.322 0.515 0.446 29.3M 153.2
Context-PIPs [36] 0.331 0.630 0.491 30.5M 146.1
EKLT [11] Event 0.254 0.842 0.174 \ \
DeepEvT [25] 0.263 0.764 0.231 185.9M 158.6
E-RAFT [12] Event 0.265 0.789 0.176 5.3M 125.4
B-FLOW [13] 0.271 0.683 0.195 5.9M 78.4
Ours Event 0.358 0.262 0.553 6.6M 239.1
Table 2: The performance of different point tracking methods on the EC and EDS datasets. The proposed approach achieves the best performance on both datasets, with particularly notable improvements on the EDS dataset, which involves more camera motion.
  Methods EDS EC
FA $\uparrow$ EFA $\uparrow$ FA $\uparrow$ EFA $\uparrow$
  EKLT [11] 0.325 0.205 0.811 0.775
DeepEvT [25] 0.576 0.472 0.825 0.818
Ours 0.616 0.529 0.854 0.834
Dataset. To validate the effectiveness of the proposed method, this paper simulates events from the PointOdyssey dataset [38], referred to here as Ev-PointOdyssey. Compared to previous datasets [9, 15], PointOdyssey offers longer durations and more annotated points on average. In practice, event generation requires a continuous visual signal, which is typically achieved by rendering high-frame-rate videos for seamless flow. Here, the method from [29] is applied to minimize pixel displacement between frames. Subsequently, DVS-voltemeter [23] is utilized to synthesize realistic event data. Some examples from the Ev-PointOdyssey dataset are shown in Fig. 4.
To evaluate performance on real-world data, experiments are also conducted on the Event Camera (EC) dataset [27] and Event-aided Direct Sparse Odometry (EDS) dataset [17]. The EC dataset includes 240×180 resolution event streams and videos recorded with a DAVIS240C sensor. The EDS dataset contains videos and events captured simultaneously using a beam splitter, with the event data recorded at 640×480 resolution by the Prophesee Gen 3.1 sensor.
Metrics. For Ev-PointOdyssey, this paper follows the experimental setup of [38], using $\sigma_{avg}$, MTE, and $Survival_{50}$ for evaluation. $\sigma_{avg}$ measures the percentage of trajectories within error thresholds of $\{1,2,4,8,16\}$, averaged across these values. Median Trajectory Error (MTE) calculates the distance between predicted and the ground truth, using the median to reduce the impact of outliers. $Survival_{50}$ indicates the average duration until tracking failure, expressed as a percentage of the total sequence length. Failure is defined as an L2 distance exceeding $50$ pixels. For fair comparison, the temporal window for TS is aligned with the ground-truth time resolution. Additionally, model parameters and inference speed are reported for comparing the resource demands of different methods.
The EC and EDS datasets, which are commonly used to evaluate event-based feature tracking methods, provide ground truth for feature points. Quantitative metrics follow the setup of [25], utilizing Feature Age (FA) and Expected Feature Age (EFA). FA measures the percentage of successful tracking steps across thresholds from $1$ to $31$, with the final score being the average across all thresholds. EFA quantifies the lost tracks by calculating the ratio of stable tracks to ground truth and scaling it by the feature age.
Implementation details. The model is trained on the Ev-PointOdyssey dataset with event clips at a spatial resolution of $256\times 320$ and a temporal length of $1.6$ seconds, optimized using Mean Absolute Error (MAE) loss. The AdamW optimizer [24] and OneCycleLR scheduler [32] are applied with a maximum learning rate of $45e-$ 4 and a cycle percentage of $0.1$. Subsequently, the model is fine-tuned with the same temporal length but a higher resolution of $512\times 640$. Experiments are conducted in parallel on 4 Nvidia RTX A6000 GPUs, implemented in PyTorch.
4.1 Quantitative Comparison
Baselines. The proposed method is compared with video-based approaches for TAP such as PIPs [15], PIPs++ [38], TAPIR [10], and Context-PIPs [36] to validate the advantages of the event modality in this task. In the absence of event-based approaches for TAP, we extend several state-of-the-art (SOTA) point correspondence methods based on events. EKLT [11] and DeepEvT [25] are event-based feature tracking methods. EKLT is built on first principles of events and can be directly applied to this dataset, while DeepEvT is retrained on the Ev-PointOdyssey dataset. E-RAFT [12] and B-FLOW [13] serve as event-based optical flow estimation methods. We link the optical flow across times, using bilinear interpolation to calculate flow for sub-pixel points. Predicted coordinates for points that exceed the boundaries are clamped to the boundary.
(a) Translation
(b) Rotation
Figure 5: The two figures present the results of dense point tracking by the proposed method on the EC dataset. (a) Camera translation sequence. (b) Camera rotation sequence.
(a) PIPs++ [38]
(b) E-RAFT [12]
(c) DeepEvT [25]
(d) Ours
(e) Ground Truth
Figure 6: Qualitative results for EDS (top two rows), Ev-PointOdyssey (bottom two rows). The predicted trajectories are visualized using a pink-to-yellow colormap, with sparse ground truth marked by “$\times$”. The first row highlights the tracking performance of different methods in low-texture regions (white tabletop), where DeepEvT almost loses tracking. In the second row, the target point (table corner) moves out of view and then returns, causing E-RAFT to fail because it only models inter-frame pixel displacement. The last two rows show dense tracking results on the Ev-PointOdyssey dataset, highlighting the superiority of the proposed method.
Ev-PointOdyssey results. On the Ev-PointOdyssey dataset, the proposed method outperforms video-based methods across all three metrics, as shown in Tab. 1. Video-based methods rely on iterative local searches to match appearance information, which leads to errors when faced with large displacements or nonlinear motion. In contrast, the proposed method incorporates kinematic features to guide matching, effectively addressing these challenges. Compared to event-based methods, EKLT and DeepEvT are designed for feature tracking as they directly predict the displacement of tracking points from local event streams. However, they struggle in low-texture areas due to insufficient event data, causing tracking failures. The proposed method integrates spatial context through local-global iterative matching, allowing tracking in low-texture regions. E-RAFT and B-FLOW estimate pixel displacement over short intervals but are susceptible to disruption from occlusions or points moving out of bounds. Although this paper does not explicitly model occlusions, its ability to capture long-term motion dependencies enables it to maintain stable tracking even when points become temporarily invisible. Moreover, the proposed method is parameter-efficient and provides faster runtime. This efficiency stems from two factors: first, event cameras focus solely on dynamic changes, thereby reducing visual redundancy; second, a transformer replaces the MLPs [38, 36] to effectively capture temporal dependencies. As EKLT is not a deep learning algorithm, a fair comparison of its parameter count and FPS is not feasible, therefore it is not included in the table.
EC and EDS results. Similar to the results on Ev-PointOdyssey, the proposed method outperforms existing event-based trackers on real-world datasets, as reported in Tab. 2. The EDS dataset, with faster camera motion than the EC dataset, results in generally lower performance across all methods. Nevertheless, the proposed method effectively handles the noise introduced by this, ensuring stable tracking results. While these metrics reflect feature tracking performance, the proposed approach also demonstrates superior performance on TAP. Figure 5 presents two sequences from the EC dataset: one with translational camera motion and the other with rotational camera motion, highlighting the robustness of the proposed method in tracking dense points across diverse motion patterns.
4.2 Qualitative Analysis
Figure 6 illustrates a comparison of the proposed method with prior works on the EDS and Ev-PointOdyssey datasets. PIPs++ takes a sequence of 48 RGB frames as input, while other methods rely on event data corresponding to the same time period. Trajectories are shown in a pink-to-yellow colormap, indicating point movement from the pink starting position to the yellow endpoint. Due to space constraints, 48 ground truth points are downsampled to 9, marked with an “$\times$” pattern in the first two rows.
The top two rows show the motion of the tabletop (a low-texture point) and the table corner (a key point). The tabletop point remains consistently visible, while the table corner temporarily exits the field of view before reentering. The first row depicts the predicted trajectory of the tabletop point across different methods. Here, DeepEvT struggles due to its reliance on local event data, which limits accuracy in low-texture areas with sparse events, leading to missed points. The second row highlights the predicted trajectory of the table corner point. E-RAFT fails to maintain tracking when points move out of the frame because it models only short-term pixel displacements rather than the motion of physical surface points. For both types of points, PIPs++ exhibits lower tracking accuracy under rapid motion compared to our method, particularly on EDS, which involves intense camera movement. The bottom two rows display dense tracking results on Ev-PointOdyssey, further demonstrating the superiority of the proposed approach.
Table 3: Ablation studies on each part of the proposed method.
       MGM VMAM Ev-PointOdyssey
PF MLP CA TC TA $\sigma_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  0.324 0.385 0.485
0.335 0.367 0.489
0.340 0.326 0.495
0.349 0.291 0.510
0.348 0.288 0.506
0.358 0.262 0.553
(a) Kinematic vectors w/o MLP
(b) Kinematic vectors w/ MLP
Figure 7: Impact of the MLP in MGM. Arrows represent the direction of kinematic vectors, and their length indicates intensity.
4.3 Ablation Study
Impact of the motion-guidance module. The first three rows of Tab. 3 illustrate how kinematic features from MGM contribute to point displacement updates, where PF stands for Plane Fitting. The first row shows a baseline model without MGM, relying solely on appearance feature matching. The second row includes MGM but omits MLP-based correction for ambiguous kinematic features. Results indicate that kinematic features play a significant role in enhancing tracking accuracy, and this effect is further amplified when corrected by the MLP. As shown in Fig. 7, the initial kinematic features are inaccurate at some time steps due to object motion overlap, but they exhibit temporal coherence after MLP correction.
Effectiveness of variable motion aware module. Rows four through six in Tab. 3 present the ablation studies for each component of VMAM, where CA, TC, and TA represent Cross Attention, Temporal Convolution, and Temporal Attention, respectively. The fourth row performs correlations solely within appearance features, without guidance from kinematic features. The fifth row employs 1D temporal convolutions to capture short-term dependencies. In the final row, the initial experimental setup is maintained, modeling temporal relationships via both short-term and long-term paths. The results reveal that incorporating both kinematic guidance and temporal modeling progressively enhances performance, highlighting the effectiveness of each VMAM component. Figure 8 provides a visual comparison. Features without VMAM exhibit temporal discontinuities, while VMAM-enhanced features demonstrate improved temporal consistency with lower standard deviation.
(a) Temporal distribution
(b) Standard deviation
Figure 8: Effectiveness of VMAM. The first row shows results using only appearance features, without VMAM, while the second row includes VMAM. (a) Temporal distribution across feature channels. (b) Temporal standard deviation across feature channels.
5 Conclusion
This study introduces a novel event-based framework for tracking any point, leveraging a motion-guidance module to extract kinematic features that refine the matching process and constructs a dynamic-appearance space. Additionally, the integration of a variable motion aware module enables the system to account for motion variations, ensuring temporal consistency across diverse velocities. To evaluate the approach, a simulated event point tracking dataset was collected. The proposed method achieved state-of-the-art tracking performance on both the simulated dataset and two real-world datasets, with competitive model parameters and faster inference time. This technique holds substantial potential for applications in embodied intelligence, autonomous driving, and related fields.